Evaluate the integral 1/2"\pi" i"\oint"dz /z2(z2+2z+3), where c is the circle |z|=3.


1
Expert's answer
2021-09-06T16:06:50-0400

"\frac{1}{2\pi i}\oint\limits_{|z|=3} \frac{dz}{z^2(z^2+2z+3)}=?"

All the roots of the denominator are: "0, 0, -1+\sqrt{2}, -1-\sqrt{2}". They all lie inside the contour "C_3=\{|z|=3\}". Thus, if "R>3", then there are no roots between the contours "C_3" and "C_R=\{|z|=R\}". Therefore, by the Cauchy integral theorem, for all "R>3",

"\frac{1}{2\pi i}\oint\limits_{|z|=3} \frac{dz}{z^2(z^2+2z+3)}=\frac{1}{2\pi i}\oint\limits_{|z|=R} \frac{dz}{z^2(z^2+2z+3)}"

and therefore

"\frac{1}{2\pi i}\oint\limits_{|z|=3} \frac{dz}{z^2(z^2+2z+3)}=\lim\limits_{R\to+\infty}\frac{1}{2\pi i}\oint\limits_{|z|=R} \frac{dz}{z^2(z^2+2z+3)}"

But

"|\frac{1}{2\pi i}\oint\limits_{|z|=R} \frac{dz}{z^2(z^2+2z+3)}|\leq \frac{1}{2\pi}\oint\limits_{|z|=R} \frac{|dz|}{|z|^2|z^2+2z+3|}\leq"

"\leq \frac{1}{2\pi}\oint\limits_{|z|=R} \frac{|dz|}{|z|^2(|z|-3)^2}= \frac{1}{2\pi}\frac{2\pi R}{R^2(R-3)^2}=\frac{1}{R(R-3)^2}\to 0"

therefore, "\frac{1}{2\pi i}\oint\limits_{|z|=3} \frac{dz}{z^2(z^2+2z+3)}=\lim\limits_{R\to+\infty}\frac{1}{2\pi i}\oint\limits_{|z|=R} \frac{dz}{z^2(z^2+2z+3)}=0".

Answer. "\frac{1}{2\pi i}\oint\limits_{|z|=3} \frac{dz}{z^2(z^2+2z+3)}=0".